# Three-Field String Inflation with Perturbative Corrections: Dynamics and Implications

### v1

@@
Vasileios Basiouris  [^1] E-mail: v.basiouris@uoi.gr, Dibya Chakraborty  [^2] E-mail: dibyac@physics.iitm.ac.in
@@
In this work, we construct an explicit string motivated example of three-field inflation in a related, yet distinct from, the recently discovered perturbative large volume scenario (pLVS). Contrary to the usual constructions, in this set up, large volume is ensured by the interplay between the effects of $\alpha^{\prime 3}$, logarithmic loop and higher derivative $F^{4}$ corrections. After addressing a full moduli stabilization scenario, we move on to a detailed analysis of three-field model of inflation after choosing a suitable canonical basis for the fields. We conduct multiple consistency checks to establish a solid foundation for our model within the framework of the underlying 4D effective field theory (EFT). Our model differs from previous setups in three key aspects: first, the interaction between subleading corrections that drive full moduli stabilization follows a different pattern, second, the volume form of the underlying Calabi-Yau is different, and third, in our three-field inflation scenario, the second slow-roll parameter consistently dominates over the first by several orders of magnitude. The latter signals the possible presence of primordial features which can be verified by forthcoming ground and space based experiments. We can roughly distinguish two stages of inflation: the first stage mostly occurs in the steepest direction during horizon crossing giving us almost $55$ efolds of inflation– once one of the inflatons falls off the ridge and then to its true minimum, the other two fields become active, giving us a truly multi-field behavior in the second stage – adding few more efodls of inflation. We also confirm our claim by introducing the non-planar torsion in the inflationary trajectory – this quantity becomes non-trivial in the second stage of inflation. Finally, we calculate the cosmological observables, which align with Planck data, and discuss potential directions for future research.
@@ 1 Introduction
Cosmic inflation refers to a rapid expansion of spacetime occured in the early Universe Guth ; Linde81 ; AS82 , which smoothed out any fluctuations in spacetime, resulting in a homogeneous and isotropic cosmos. Our theoretical understanding of the Universe and its evolution is based on the $\Lambda$CDM model, a six-parameter framework that provides a remarkably accurate explanation for most of the cosmological phenomena. However, it lacks an explanation for the horizon problem and the observed flatness of our Universe. Inflation can provide an explanation for the above two, besides, it can also give rise to the seed for the structure formation we observe today. In addition, observations such as Planck Planck:2018jri are fully consistent with the simplest inflationary scenario being the leading mechanism to account for the observed anisotropies in the Cosmic Microwave Background (CMB) radiation. Specifically, they align with two strong predictions of inflation: a nearly scale-invariant power spectrum of density perturbations and a stochastic background of gravitational waves.
According to the simplest inflationary scenario, inflation is driven by a scalar field minimally coupled to gravity, whose potential energy dominates over its kinetic energy, causing it to slowly roll over its potential. Once the kinetic energy begins to dominate, inflation comes to an end, providing sufficient spacetime expansion—measured by a parameter called e-folds to resolve the horizon problem. Hence, for inflation to come to a stop eventually, we require a quasi de Sitter acclerated expansion. However, recently proposed consistentcy conjectures Vafa:2005ui ; Obied:2018sgi ; Agrawal:2018own ; Palti:2019pca imposed on the 4D effective field theory (EFT) derived from a quantum theory of gravity, strongly disfavors vanilla single field inflation. Although, it was hinted upon in Achucarro:2018vey that multi-field inflation with non-trivial field space metric might be able to alleviate this tension.
@@
Four-dimensional EFT emerging after dimensional reduction of a superstring theory features various scalar fields, and one/some of them might act as natural candidate/s to drive inflation. This forms the foundation of our three-field inflation model. However, these constructions also have a price to pay; most of the times, the existence of too many scalar fields becomes unavoidable, leading to the so-called cosmological moduli problem (see Cicoli:2023opf and the references therein). One also needs to stabilize the scalar field to prevent the appearance of fifth forces. Our aim in this paper is to display a concrete three-field inflationary scenario where the potential energy is constructed from an underlying string theory setup, in particular from a type-IIB flux compactification where the extra dimensions are compactified on a K3 fibred Calabi-Yau (CY). Hence, the first step towards string theory model building is to stabilize the scalar fields. Fortunately, string theory provides sufficient tools (such as $Dp$ branes pierced with magnetized fluxes) to address these issues.
@@
The scalar fields stemming from the closed string sector of type-IIB flux compactifications are of three types: complex-structure $(U^{i})$ related to the shapes of the CY, Kähler moduli $(T_{\alpha})$ related to the size of the CY, and an Universal scalar field known as axio-dilaton $(s)$ whose vacuum expectation value (vev) determines the string coupling $g_{s}$. After dimensional reduction, the scalar potential can be expressed in terms of a holomorphic superpotential and a real Kähler potential. The Type-IIB superpotential depends on the complex structure moduli and dilaton, however it is independent of the Kähler moduli, resulting in the so-called no-scale structure. Hence, to stabilize the Kähler moduli, one has to go beyond and add further quantum corrections, such as perturbative corrections to the Kähler potential and non-perturbative effects at the level of superpotential. Note that, the complex-structure and axio-dilaton are usually stabilized at tree-level in presence of fluxes and integrated out.
@@
The most notable candidate models, considering the non-perturbative nature of corrections, are the KKLT Kachru:2003aw ; Polchinski:2015bea ; Gautason:2019jwq ; Hamada:2019ack ; Linde:2020mdk ; Randall:2019ent ; Blumenhagen:2019qcg ; Andriot:2020wpp ; Abe:2005rx ; Carta:2019rhx model and the Large Volume Scenario (LVS) Conlon:2005ki ; Balasubramanian:2005zx ; Burgess:2003ic ; Cremades:2007ig . Both type of models address stabilization of the Kähler moduli in type IIB flux compactification despite being flat direction at tree-level. The former utilises non-perturbative effects to be added in a holomorphic superpotential of the Gukov-Vafa-Witten form. The latter addresses the moduli stabilization by a careful interplay between perturbative effects at the level of real Kähler potential and non-perturbative effects to the superpotential. These balance helps to achieve an exponentially large volume of the underlying 6D Calabi-Yau manifold spanned by the Kähler moduli.
@@
KKLT and LVS being the main two aspects to address full moduli stabilization, recently another class of string vacua also emerged Antoniadis:2019rkh ; Antoniadis:2018hqy ; Basiouris:2020jgp ; Basiouris:2021sdf ; Leontaris:2022rzj where it was shown that large volume can be achieved even without the need of non-perturbative effects as well as to provide an alternative explanation to the long standing problem of Dine-Seiberg Dine:1985he – this type of construction is known as pLVS. It is well known that non-perturbative effects are very restrictive and very specific towards the underlying CY geometry, addition of them may lead to satisfying several constraints Witten:1996bn . Therefore, if a construction ensures a large internal volume without relying on these effects, it becomes much more well-motivated. A recent study Bera:2024zsk provided a more detailed analysis of moduli stabilization and the potential for inflation within this pLVS framework. It also examined the robustness of pLVS against various subleading corrections, including leading-order corrections in inverse string tension, higher-derivative corrections, log-loop corrections from intersecting space-filling $D7$ brane stacks, D-term effects, and Kaluza-Klein as well as winding-type loop corrections.
Supported by these facts, the present work studies the quantum loop corrections in the context of fibre inflation like models Cicoli:2008gp ; Cicoli:2016xae ; Cicoli:2017axo , which are promising for describing inflation. Their advantages have to do with the fact that they include a larger number of cycles $h^{1,1}>3$ allowing for non-trivial $D7$ branes fluxes to introduce chiral matter to the model. Moreover, fibre inflationary paradigm can be identified as a solution to the $\eta$ problem in inflation, upgrading the fibre inflation models as ideal candidate to study the expansion of spacetime. In this lore, we are going to present a stabilization scheme, without introducing any non-perturbative effects, where the log-loop corrections together with higher derivative corrections are sufficient to stabilize the "Swiss-cheese" like Calabi-Yau. Furthermore, we explicitly compute on the transformation between the moduli and the canonically-normalized fields, where approximated solutions are found and consistency bounds for the free parameters of the theory are imposed.
Similar to studying any model of inflation, we aim to start by laying the framework to derive the scalar potential in terms of the Kähler moduli. Due to the presence of three-fields, we also discuss the non-planar trajectories and their implications. The effective expansion of spacetime occurs in two stages: initially along the steep direction of the potential, followed by expansion along the remaining two directions as the fields move toward their minima. Note that, throughout inflation all the three fields are active, but their velocity changes while swiftly transitioning between these two stages. In the later phase, the field that fell into the ridge moves relatively slowly, whereas, in the earlier stage, the other two fields move slowly. Since the fields are canonical, the non-planar motion and deviation from a geodesic occur only when the field temporarily falls into a ridge and oscillates there. Finally, we discuss the inflationary predictions and conclude the possible outlook.
@@ 2 Towards moduli stabilization with perturbative effects
The perturbative string loop corrections has been recently highlighted in Antoniadis:2018hqy due to their major role in the moduli stabilization mechanism without the need for a non-perturbative effects like the vanilla Large Volume Scenario Cicoli:2008gp . In this section, we start by summarizing the world-sheet corrections which are proportional to the Euler characteristic $\chi$ of the compactification, followed by an analysis of the origin of string loop corrections within the geometric framework of $D7$ branes in type IIB string theory. After presenting a detailed derivation of the scalar potential, we will showcase several examples of dS vacua where the correct mass hierarchy remains preserved, along with the underlying supergravity approximations.
@@
$\xi=-\dfrac{\hat{\xi}(3)}{4(2\pi)^{3}}\chi\leavevmode\nobreak\ .$ (2.2)
@@
From the above, it is observed that an additional localized Einstein-Hilbert (EH) is induced in four dimensions with a coefficient proportional to $\chi$. These terms correspond to localized gravity vertices in the bulk (in particular at points where $\chi\neq 0$), where their effect is to emit Kaluza-Klein states and gravitons. We assume a background with three intersecting $D7$ branes, where we expect interactions between the closed string modes and KK states in the bulk space. These interactions exhibit in general a logarithmic behavior in the large volume limit transverse to the $D7$-branes. The new contributions to the $R^{4}$ term which are going to be moduli dependent, can be summarized by the following formula Basiouris:2020jgp :
@@
where $\phi$ is the universal modulus known as dilaton, related to string coupling as $g_{s}=e^{\langle\phi\rangle}$; $u^{i}$ is the saxion and $v^{i}$ is the axion of the complex-structure modulus, and $\tau_{\alpha}$’s are the Einstein frame four-cycle volume moduli defined as, $\tau_{\alpha}=\partial_{t^{\alpha}}\mathcal{V}$, $\mathcal{V}$ is the overall volume of the Calabi-Yau manifold expressed as $\mathcal{V}=\dfrac{1}{3!}\kappa_{\alpha\beta\gamma}t^{\alpha}t^{\beta}t^{\gamma}$. Finally, $c_{0}$ and $c_{\alpha}$’s are the universal RR axion and RR four-form axions, respectively.
@@ 2.1 The geometrical setup
Taking into account the above discussions, we now turn to the derivation of the scalar potential. Fibre inflation model, proposed more than a decade ago in Cicoli:2008gp are promising with respect to the embedding of inflation in string compactifications, since they provide a natural solution to the $\eta-$problem Cicoli:2011zz . Different variations of fibre inflation models have been proposed, where combinations of string loop corrections, Kaluza-Klein and winding loops are considered, leading into inflationary potentials with a sufficient plateau-like form.
In all of these models, the presence of non-perturbative effects to stabilize the blow-up model and the overall volume were common. Recently, another aspect of fibre inflation is also proposed where the interplay of perturbative corrections assured a large volume Bera:2024ihl as well as full moduli stabilization. The absence of non-perturbative effects eliminates the need for an exceptional del Pezzo divisor. The novelty of this setup is that the usual drawback of fibre inflation where the field space excursion was bounded by Kähler cone conditions were alleviated. In this paper, we follow a similar approach as Bera:2024ihl but consider a different form of the volume, resembling that of K3-fibred Calabi-Yau models like fibre inflation. However, as will become clear in the following subsections, our inflationary predictions deviate significantly from those of the fiber model. To proceed in our analysis, the compactified space volume is given by:
@@
2.2 The effective scalar potential and its stabilization
Having this volume-form as a starting point, our analysis will be focused on the addition of BBHL term and logarithmic corrections in the Kähler potential without considering any non-perturbative corrections. The large volume in this model is ensured by a cumulative effect of all the perturbative corrections. The Kähler potential upon the insertion of the these corrections can be written as:
@@ 2.2 The effective scalar potential and its stabilization
where $\tau_{i}$ stand for the four-cycle moduli, related to the volume of the CY as $\mathcal{V}\sim\tau^{3/2}$, $\xi$ denotes the $\alpha^{\prime}$ correction-term and the logarithmic corrections $\sim\hat{\eta}\log(..)$ are induced along each four-cycle volume moduli. One important last observation we have to make is the fact that the stabilization will completely depend on perturbative corrections. The non-perturbative corrections, which will impose stringent limitation in our construction, could be prevented by considering certain world-volume fluxes that lift fermionic zero modes leading to zero contribution from these corrections Green:1997di ; Bianchi:2011qh ; Cvetic:2012ts ; Blumenhagen:2007sm . Taking this into account, in this work, we will explore the case where $\tau_{1}$ no longer remains at small values, but it is stabilized to large values but never taking us away from the domain of large volume.Despite the fact that this approach diverges from the typical paradigm of fibred models, our proposal focuses on the implications on the logarithmic correction on the geometries like (2.9).
Regarding the complex structure moduli $z_{i}$ as well as the axion-dilaton, they are stabilized at high energies by the supersymmetric flatness conditions:
$D_{U_{i}}W=0=D_{\bar{U}_{i}}\bar{W},\;D_{s}W=0=D_{\bar{s}}\bar{W}\leavevmode\nobreak\ .$ (2.11)
This novel procedure results into one flat direction that appears at the F-term potential level as a consequence of the higher order terms scaling as $\mathcal{O}(1/\mathcal{V}^{4})$. Since the scalar potential is parametrized by the three moduli ($\tau_{1},\tau_{7},\tau_{6}$), we would like to trade one of them in order to introduce the volume $\mathcal{V}$ in the computations. This reparametrization helps us to perform the large volume expansion. So, in order to include the overall volume in our computations, we solve the equation (2.9) with respect to one modulus e.g. $\tau_{6}$,
@@
As discussed in Cicoli:2017axo ; Cicoli:2008gp , it is a unique feature of these models to preserve one flat direction. In Cicoli:2008gp ; Cicoli:2017axo , their approach is to stabilize first the internal volume and the blow-up divisor with an interplay of perturbative and non-perturbative effects, while the $\tau_{7}$ remains flat. Our approach differs from the aforementioned due to the fact that the logarithmic corrections will be made sufficient to perform the stabilization while keeping track of the remaining flat direction. The previous studies of this kind of approach requires only keeping terms that scales as $\sim\frac{1}{\mathcal{V}^{3}}$ , but now we will prove that even terms proportional to $\sim\frac{\hat{\eta}\hat{\hat{\xi}}}{\mathcal{V}^{4}}$ could be significant. Firstly by inspecting the effective potential (2.19), we stabilize with respect to to $\mathcal{V}$ while considering $\mathcal{V}\gg\tau_{1}$:
@@
where we can see that $\tau_{1,7}$ directions still remain flat. To remedy this, we now include next to leading order term:
@@
$K_{ij}\partial T_{i}\partial\bar{T_{j}}=\sum_{i}\dfrac{\partial T_{i}\partial\bar{T_{i}}}{(T_{i}+\bar{T}_{i})^{2}}=\sum_{i}\dfrac{1}{2}(\partial\varphi_{i})^{2}+...\leavevmode\nobreak\ .$ (2.34)
The new relation between the old and new bases are the following:
@@
To prove the validity of our approximate formula of the potential in (2.40) and the exact potential in (2.37), we are going to sketch the potentials in the vicinity of the global minimum.
@@
Table 1: Three different numerical solutions for the minima with their corresponding masses.
@@
Table 2: The free parameters used for the examples presented above.

### v2

@@
Vasileios Basiouris  [^1] E-mail: v.basiouris@uoi.gr, v.basiouris@gmail.com, Dibya Chakraborty  [^2] E-mail: dibyac@physics.iitm.ac.in
@@
In this work, we construct an explicit string motivated example of three-field inflation in a related, yet distinct from the recently discovered perturbative large volume scenario (pLVS). In this set up, large volume is ensured by the interplay between the effects of $\alpha^{\prime 3}$, logarithmic loop and higher derivative $F^{4}$ corrections. After addressing full moduli stabilization, we move on to a detailed analysis of a three-field model of inflation while choosing a suitable canonical basis for the fields. Our model differs from previous setups in three key aspects: first, the interaction between subleading corrections that drive full moduli stabilization is different compared to pLVS for instance; second, the volume of the underlying Calabi-Yau is of "weak-Swiss-cheese" type; and third, in our three-field inflationary scenario, the second Hubble slow-roll parameter consistently dominates over the first by several orders of magnitude — primarily due to the steepness of the potential. When this hierarchy is aided by a brief deviation from slow-roll, an enhancement in the primordial power spectrum can be naturally achieved. We have obtained two stages of inflation: the first stage mostly occurs when one of the inflatons roll along the steepest direction —completing $55$ efolds of inflation. Once the rolling inflaton falls off a ridge of the multi-field potential, the other two fields become active, giving us a true multi-field behavior in the second stage – adding few more efolds of inflation. Inflation ends when all three of them roll towards the global minimum and settles there. We confirm this behavior by introducing the non-planar torsion in the inflationary trajectory – this quantity becomes non-trivial in the second stage of inflation. Finally, we calculate the cosmological observables, which align with the Planck data, and discuss potential directions for future research.
@@ 1 Introduction
Cosmic inflation refers to a rapid expansion of spacetime occured in the early Universe Guth ; Linde81 ; AS82 , which smoothed out any fluctuations in spacetime, resulting in a homogeneous and isotropic cosmos. Our current understanding of the Universe and its evolution is based on the six-parameter $\Lambda$CDM model. The model lacks an explanation for the horizon problem and the observed flatness of our Universe. Inflation can provide an explanation for the above two, besides, it can also give rise to the seed for the structure formation we observe today. In addition, the data from Planck Planck:2018jri are fully consistent with the simplest inflationary scenario being the leading mechanism to account for the observed anisotropies in the Cosmic Microwave Background (CMB) radiation. Specifically, they align with two strong predictions of inflation: a nearly scale-invariant power spectrum of density perturbations and a stochastic background of gravitational waves.
According to the simplest inflationary scenario, inflation is driven by a scalar field minimally coupled to gravity, whose potential energy dominates over its kinetic energy, causing it to slowly roll over its potential. Once the kinetic energy begins to dominate, inflation comes to an end, providing sufficient spacetime expansion to resolve the horizon problem. However, recently proposed consistentcy conjectures Vafa:2005ui ; Obied:2018sgi ; Agrawal:2018own ; Garg:2018reu ; Palti:2019pca imposed on the 4D effective field theory (EFT) derived from a quantum theory of gravity, strongly disfavors vanilla single field inflation. Although, it was hinted upon in Achucarro:2018vey that multi-field inflation with non-trivial field space metric might be able to alleviate this tension.
@@
Four-dimensional EFT emerging after dimensional reduction of a superstring theory features various scalar fields, and one/some of them might act as natural candidate/s to drive inflation. This forms the foundation of our three-field inflation model. However, these constructions also have a price to pay; most of the time, the existence of too many scalar fields becomes unavoidable, leading to the so-called cosmological moduli problem (see Cicoli:2023opf and the references therein). One also needs to stabilize the scalar fields to prevent the appearance of fifth forces. Our aim in this paper is to display a concrete three-field inflationary scenario where the scalar potential is constructed out of the underlying low-energy effective string theory setup, in particular from a type-IIB flux compactification where the extra dimensions are compactified on a K3 fibred Calabi-Yau (CY). The first step towards string theory model building is to stabilize the scalar fields. Fortunately, string theory provides sufficient tools (such as $Dp$ branes pierced with magnetized fluxes) to help achieve this goal.
@@
The scalar fields stemming from the closed string sector of type-IIB flux compactifications are of three types: complex-structure $(U^{i})$ related to the shapes of the CY, Kähler moduli $(T_{\alpha})$ related to the size of the CY, and an Universal scalar field known as axio-dilaton $(s)$. Vacuum expectation value (vev) of the imaginary part of axio-dilaton determines the string coupling $g_{s}$. After dimensional reduction, in 4D supergravity, the scalar potential can be expressed in terms of a holomorphic superpotential and a real Kähler potential. The Type-IIB superpotential depends on the complex structure moduli and dilaton, however it is independent of the Kähler moduli, resulting in the so-called no-scale structure. Hence, to stabilize the Kähler moduli, one has to go beyond and add further quantum corrections, such as perturbative corrections to the Kähler potential and non-perturbative effects at the level of superpotential. Note that the complex structure and axio-dilaton are stabilized at tree level in the presence of fluxes and integrated out.
@@
The most notable candidate models — addressing full moduli stabilization are the KKLT Kachru:2003aw ; Polchinski:2015bea ; Gautason:2019jwq ; Hamada:2019ack ; Linde:2020mdk ; Randall:2019ent ; Blumenhagen:2019qcg ; Andriot:2020wpp ; Abe:2005rx ; Carta:2019rhx model and the Large Volume Scenario (LVS) Conlon:2005ki ; Balasubramanian:2005zx ; Burgess:2003ic ; Cremades:2007ig . The former utilises non-perturbative effects to be added in a holomorphic superpotential of the Gukov-Vafa-Witten form. The latter addresses the moduli stabilization by a careful interplay between perturbative effects at the level of real Kähler potential and non-perturbative effects to the superpotential. These balance in LVS setup helps to achieve an exponentially large volume of the underlying 6D Calabi-Yau manifold spanned by the Kähler moduli.
@@
KKLT and LVS being the main two examples to address full moduli stabilization, recently another class of string vacua also emerged Antoniadis:2019rkh ; Antoniadis:2018hqy ; Basiouris:2020jgp ; Basiouris:2021sdf ; Leontaris:2022rzj where it was shown that exponentially large volume can be achieved without the need of non-perturbative effects — this type of construction is known as pLVS. Non-perturbative effects are very restrictive and very specific towards the underlying CY geometry, additionally they may need to satisfy several constraints such as Witten’s unit arithmetic genus condition Witten:1996bn . Therefore, if a construction ensures a large internal volume without relying on these effects, it becomes more well-motivated. A recent study Bera:2024zsk have provided a global embedding of the pLVS framework and the robustness of the same against various subleading corrections, including leading-order corrections in inverse string tension, higher-derivative corrections, log-loop corrections from intersecting space-filling $D7$ brane stacks, D-term effects, and Kaluza-Klein as well as winding-type loop corrections.
The present work studies the addition of perturbative effects such as leading $\alpha^{\prime 3}$ correction (BBHL term), corrections of log-loop type and from higher derivative terms, in a weak Swiss-cheese CY. We start deriving the multi-field scalar potential and then disucss the stabilization of the scalars putting special emphasis on the higher derivative $F^{4}$ effect and the log-loop corrections. Then we move on to discuss their inflationary implications. Inflation takes place in two stages: initially along the steep direction of the potential, and then along the remaining two directions as all the fields move towards their global minimum from an initially displaced position. At this initially displaced position in the scalar potential, the minimum along one of the scalars disappears, producing a steep ridge in that direction. The inflaton then falls in the ridge and oscillates — in this stage the other two fields move relatively slower, giving rise to a quassi-single field behavior in the first stage. In the later stage, the field that fell into the ridge moves relatively slowly. Since the fields are canonical, the non-planar motion and deviation from a geodesic occur only when the field temporarily falls into a ridge and oscillates there. The oscillation can not only give rise to high turning rate and torsion, but it can also induce sizable entropic perturbation through transient deviation from slow-roll. The backreaction of the entropic modes to the adiabatic modes may produce a significant enhancement in the primordial power spectrum signalling a copious production of primordial blackholes (PBH). In this lore, it is also useful to study the effect of slow-roll violation in the secondary induced gravitational waves. We aim to come back to the implications of slow-roll violation in a future companion paper.
@@ 2 Towards moduli stabilization with perturbative effects
The perturbative string loop corrections has been recently highlighted in Antoniadis:2018hqy due to their major role in the moduli stabilization mechanism without the need for a non-perturbative effects like the Large Volume Scenario Cicoli:2008gp . In this section, we start by summarizing the world-sheet corrections which are proportional to the Euler characteristic $\chi$ of the compactification, followed by an analysis of the origin of string loop corrections within the geometric framework of $D7$ branes in type IIB string theory. After presenting a detailed derivation of the scalar potential, we shall showcase several examples of dS vacua where the correct mass hierarchy remains preserved.
@@
$\xi=-\dfrac{\zeta(3)}{4(2\pi)^{3}}\chi\leavevmode\nobreak\ .$ (2.2)
@@
From the above, it is observed that an additional localized Einstein-Hilbert (EH) is induced in four dimensions with a coefficient proportional to $\chi$. These terms correspond to localized gravity vertices in the bulk (in particular at points where $\chi\neq 0$), where their effect is to emit Kaluza-Klein states and gravitons. We assume a background with three intersecting $D7$ branes, where we expect interactions between the closed string modes and KK states in the bulk space. These interactions exhibit in general a logarithmic behavior in the large volume limit transverse to the $D7$-branes. The new contributions to the ${R}_{(4)}$ term which are going to be moduli dependent, can be summarized by the following formula Basiouris:2020jgp :
@@
where $\phi$ is the universal modulus known as dilaton, related to string coupling as $g_{s}=e^{\langle\phi\rangle}$; $u^{i}$ is the saxion and $v^{i}$ is the axion of the complex-structure modulus, and $\tau_{\alpha}$’s are the Einstein frame four-cycle volume moduli defined as, $\tau_{\alpha}=\partial_{t^{\alpha}}\mathcal{V}$, $\mathcal{V}$ is the overall volume of the CY manifold in Einstein frame expressed as $\mathcal{V}=\dfrac{1}{3!}\kappa_{\alpha\beta\gamma}t^{\alpha}t^{\beta}t^{\gamma}$. Finally, $c_{0}$ and $c_{\alpha}$’s are the universal RR axion and RR four-form axions, respectively.
@@ 2.1 The geometrical setup
Taking into account the above discussions, we now turn to the derivation of the scalar potential and discuss the CY geometry. Fibre inflation model, proposed more than a decade ago in Cicoli:2008gp is promising with respect to the embedding of inflation in string compactifications, since they provide a natural solution to the $\eta-$problem of inflation Cicoli:2011zz . Different variations of fibre inflation models have been proposed, where combinations of string loop corrections, Kaluza-Klein and winding loops are considered, leading into inflationary potentials with a sufficient plateau-like form.
In all of these models, the presence of non-perturbative effects to stabilize the blow-up model and the overall volume were common. Recently, another aspect of fibre inflation is also proposed where the interplay of perturbative corrections assured a large volume Bera:2024ihl as well as full moduli stabilization. The absence of non-perturbative effects eliminates the need for an exceptional del Pezzo divisor. The novelty of this setup is that the usual drawback of fibre inflation where the field space excursion was bounded by Kähler cone conditions were alleviated. The use of perturbative corrections in the context of inflationary models can also be found in Bansal:2024uzr where the perturbative effects were used in the stabilization of the blow-up mode. Later, from the perspective of brane-antibrane inflation, these effects are also proven useful in Cicoli:2024bwq .
In this paper, we follow a similar approach as Bera:2024ihl but consider a different form of the volume, resembling that of K3-fibred Calabi-Yau models like fibre inflation. However, as will become clear in the following subsections, our inflationary predictions deviate significantly from those of the fibre models because we do not utlize non-perturbative effects to stabilize the overall volume and the blow-up modulus. To proceed with our analysis, the compactified space volume is given by:
@@
2.2 The effective scalar potential and moduli stabilization
Having this volume-form as a starting point, our analysis will be focused on the addition of BBHL term and logarithmic corrections in the Kähler potential without considering any non-perturbative corrections. The large volume in this model is ensured by a cumulative effect of the perturbative corrections. The Kähler potential upon the insertion of the these corrections can be written as:
@@ 2.2 The effective scalar potential and moduli stabilization
where $\tau_{i}$ stand for the four-cycle moduli, related to the volume of the CY as $\mathcal{V}\sim\tau^{3/2}$, $\xi$ denotes the $\alpha^{\prime}$ correction-term and the logarithmic corrections $\sim\hat{\eta}\log(..)$ are induced along each four-cycle volume moduli. One important last observation we have to make is the fact that the stabilization will completely depend on perturbative corrections. The non-perturbative corrections, which will impose stringent limitation in our construction, could be prevented by considering certain world-volume fluxes that lift fermionic zero modes leading to zero contribution from these corrections Green:1997di ; Bianchi:2011qh ; Cvetic:2012ts ; Blumenhagen:2007sm . Taking this into account, in this work, we will explore the case where $\tau_{1}$ no longer remains at small values, but it is stabilized to large values but never taking us away from the domain of large volume.Despite the fact that this approach diverges from the typical paradigm of fibred models, our proposal focuses on the implications on the logarithmic correction on the weak-Swiss-cheese CY geometry of (2.9).
Regarding the complex structure moduli $U^{i}$ as well as the axion-dilaton, they are stabilized at tree-level by the supersymmetric conditions:
$D_{U^{i}}W=0=D_{\bar{U}^{i}}\bar{W},\;D_{s}W=0=D_{\bar{s}}\bar{W}\leavevmode\nobreak\ .$ (2.11)
We will also set:
$\mathcal{W}_{0}=\left\langle\int_{X_{6}}G_{3}\wedge\Omega\right\rangle$ (2.12)
where $G_{3}=F_{3}+i\,sH_{3}$ with $F$ and $H$ denote RR and NSNS 3-form fluxes respectively. $\Omega$ is the holomorphic $(3,0)-$form which depends on the complex sturcture moduli $U^{i}$. Since the scalar potential is parametrized by the three moduli ($\tau_{1},\tau_{7},\tau_{6}$), we would like to trade one of them in order to introduce the volume $\mathcal{V}$ in the computations. This reparametrization helps us to perform the large volume expansion. So, in order to include the overall volume in our computations, we solve the equation (2.9) with respect to one modulus e.g. $\tau_{6}$,
@@
As discussed in Cicoli:2017axo ; Cicoli:2008gp , it is a unique feature of these models to preserve one flat direction. In Cicoli:2008gp ; Cicoli:2017axo , their approach is to stabilize first the internal volume and the blow-up divisor with an interplay of perturbative and non-perturbative effects, while the $\tau_{7}$ remains flat. Our approach differs from the aforementioned due to the fact that the logarithmic corrections will be made sufficient to perform the stabilization while keeping track of the remaining flat direction. The previous studies of this kind of approach requires only keeping terms that scales as $\sim\frac{1}{\mathcal{V}^{3}}$ , but now we shall prove that even terms proportional to $\sim\frac{\hat{\eta}\hat{\hat{\xi}}}{\mathcal{V}^{4}}$ could be significant. Firstly by inspecting the effective potential (2.20), we stabilize with respect to to $\mathcal{V}$ while considering $\mathcal{V}\gg\tau_{1}$:
@@
where we can see that $\tau_{1,7}$ directions still remain flat. To remedy this, we now include the next to leading order term:
@@
$K_{ij}\partial T_{i}\partial\bar{T_{j}}=\sum_{i}\dfrac{1}{2}(\partial\varphi_{i})^{2}+...\leavevmode\nobreak\ .$ (2.35)
The new relation between the old and new bases where the off-diagonal entries are found small in the $\varphi\sim\mathcal{O}(1)$ limit and ignored, are the following:
@@
To prove the validity of our approximate formula of the potential in (2.41) and the exact potential in (2.38), we are going to sketch the potentials in the vicinity of the analytical and numerical global minimum.
@@
Table 1: Three different numerical solutions and the corresponding mass scales.
@@
Table 2: The free parameters used for the examples presented in table 1.